# Recent Advances in Named Entity Recognition: A Comprehensive Survey and Comparative Study

### v2

@@ 5.3 Deep-learning-based methods
The introduction of neural probabilistic language models by Bengio et al. (2003) laid the foundation for deep learning in NLP. By demonstrating the power of distributed word representations, Bengio’s work showed that neural networks could capture word similarities in high-dimensional spaces, a capability that would later be crucial for deep learning-based NER systems. Building on this foundation, Collobert (2011) applied deep learning to NER, proving that CNNs could be successfully used for a wide range of NLP tasks, including NER, semantic role labeling, and chunking.
@@ 5.3.1 Data representation
Word embeddings transform words into dense (or sparse) vectors of real numbers in a high-dimensional space, potentially up to 4000 dimensions. The resulting vectors represent each word’s semantic meaning based on surrounding words in a corpus. Traditional methods include techniques such as One-Hot encoding and TF-IDF (Ramos et al., 2003). Static word embeddings, such as Word2Vec (Mikolov et al., 2013a), GloVe (Pennington et al., 2014), and fastText (Bojanowski et al., 2016), provide a fixed vector for each word regardless of its context. Contextual embeddings, based on transformers such as GPT (Radford et al., 2018) and BERT (Devlin et al., 2018a), capture the meaning of words based on their specific context within a sentence, allowing for more nuanced and flexible word representations.
@@
Modern word embeddings like Word2Vec and GloVe produce dense vector representations, where similar terms have close vectors. These embeddings are derived from large corpora in an unsupervised manner. Word2Vec uses Continuous Bag-of-Words (CBOW), which predicts a target word from its context, and Skip-gram, which predicts context words from a target word. Both architectures are trained with noise contrastive estimation (NCE), a type of negative sampling. NCE increases the likelihood of the target word in context and decreases the likelihood of noise words, effectively teaching the model to distinguish between true contexts and artificially generated noise (Mikolov et al., 2013b). This contrastive approach leads to dense embeddings that are highly effective for capturing semantic similarities, as it forces the model to focus on discriminative features during training, enhancing the quality of the resulting vector representations. GloVe, on the other hand, uses matrix factorization of the word co-occurrence matrix, capturing the corpus’s general statistics for efficient, high-performance embeddings (Pennington et al., 2014). Unlike Word2Vec, which is context-based, GloVe emphasizes global word co-occurrence to produce vectors that represent the global relationships between words.
Numerous studies, including (Collobert et al., 2011; Huang et al., 2015), have used word embeddings for NER. For example, Ma and Hovy (2016) evaluated the performance of their NER system using different word embeddings such as Word2Vec and GloVe. The importance of these embeddings for achieving good performance is presented in (Lample et al., 2016).
Finally, we note that word embeddings can also be combined, as in (Dadas, 2019; Das et al., 2017), where a Wikipedia knowledge base is used to annotate named entities. In (Dadas, 2019), the labels are transformed into One-Hot vectors and concatenated with Word2Vec or ELMo (Peters et al., 2018a) word embeddings.
Character embeddings represent words as sequences of character vectors, capturing their internal structure. Unlike traditional embeddings, they vectorize characters and combine them to form word representations, as Figure 6 illustrated in the case of a CNN architecture.
@@
Character embeddings can be created using several methods. One-hot encoding is a technique that represents characters as binary vectors with a 1 at the index corresponding to the character. Chars2Vec [^1] https://github.com/IntuitionEngineeringTeam/chars2vec?tab=readme-ov-file, inspired by Word2Vec, generates embeddings by predicting characters based on their neighboring characters. CNN-based embeddings treat character sequences as images, using CNN filters to extract meaningful features. BiLSTM-based embeddings use bidirectional Long Short-Term Memory networks to capture contextual information from character sequences.
CNN and RNN, which we describe in the following section, are used in (Ma and Hovy, 2016; Peters et al., 2017, 2018b; Zhou et al., 2023a) to calculate feature vectors for each word. One of the conclusions drawn in (Lample et al., 2016) is that recurrent models tend to prioritize later elements, leading to feature vectors that represent suffixes more strongly than prefixes. Consequently, the authors suggest the use of BiLSTM to more effectively capture the prefix information.
@@ 5.3.2 Context encoding
It is for this reason that Huang et al. (2015) introduced an LSTM model for NER and showed that adding a CRF layer as a tag decoder can improve performance. This combination of LSTM for sequence modeling and CRF for decoding proved to be a highly effective strategy. In other areas, similar approaches were used by Chalapathy et al. (2016) for Drug NER and Zhang and Yang (2018) for Chinese NER. An RNN based on a BiLSTM framework was used in (Huang et al., 2015); the same approach has since been adopted by other authors (Ma and Hovy, 2016; Lample et al., 2016; Huang et al., 2020a). This widespread adoption underscores the robustness and flexibility of RNN-based models for a variety of NER applications.
@@ 5.3.3 Entity decoding
CRF, for example, is a probabilistic model that assigns labels by considering the entire sequence of tokens rather than making independent predictions for each one. It excels in modeling dependencies between adjacent labels, which is crucial for ensuring coherent tag sequences. By learning to assign higher probabilities to valid label sequences (such as recognizing that a "Person" label is likely to be followed by a "Location"), CRF provides a robust method for NER. Studies by Lample et al. (2016) and Ma and Hovy (2016) show that incorporating a CRF layer into deep learning models, such as BiLSTM (Luo et al., 2018; Lin et al., 2019) or CNN (Knobelreiter et al., 2017; Feng et al., 2020), enhances performance. This combination leverages CRF’s ability to ensure consistent sequence labeling while benefiting from the feature extraction power of neural networks. Many advanced NER models integrate a CRF layer with either BiLSTM or CNN for feature extraction. The BiLSTM captures the sequential context by processing the input in both forward and backward directions, ensuring that the model considers the entire sentence before making predictions. CNN, on the other hand, focuses on local patterns such as prefixes, suffixes, and word combinations, which are key in identifying entities. The CRF layer is then used to refine the final output, ensuring consistency across the sequence. This combination allows the model to capture both local features and long-range dependencies, making it more accurate and reliable.
@@ 5.4 Transformer-based language methods
Although Transformer-based models such as BERT, RoBERTa, DeBERTa, and GLiNER have shown excellent performances in NER, they are computationally intensive. The motivation behind attempts to create more efficient models such as DistilBERT is obtaining a better balance between performance and resource requirements. Recent advances also include the development of multilingual models such as mBERT (Devlin et al., 2018b), which can handle multiple languages and be fine-tuned for specific languages using bilingual gazetteers or additional datasets, enhancing performance in low-resource settings. Additionally, models such as Language Understanding with Knowledge-based Embeddings (LUKE) incorporate entity information into word embeddings, improving entity recognition by using innovative masking and self-attention techniques (Yamada et al., 2020).
@@ 5.5.1 Principles and applications of LLMs in NER
LLMs rest on the Transformer decoder architecture, in which a multitude of attention mechanisms are orchestrated in layers to form an intricate neural network. The structural designs and pre-training paradigms implemented in current LLMs exhibit strong parallels with those used in smaller-scale language models. The primary distinction is the markedly increased size of both the model parameters and the training corpus. Some LLMs, such as T5 (Raffel et al., 2020), operate as hybrids, incorporating both the encoder and decoder modules of the Transformer to increase comprehension and generative functionalities.
@@ 5.5.2 Pros and Cons of LLMs for NER
Another important advantage is contextual understanding. The transformer-based architecture of LLMs enables these models to capture complex contextual relationships, an essential feature for NER tasks in specialized domains, such as medical or historical texts. For instance, models like Resolving ECCNPs (Kammer et al., 2023), tailored for German medical texts, effectively handle unique language constructs and domain-specific terminology. This ability to understand nuanced context improves LLM performance in scenarios where accurate interpretation of specialized language is crucial.
@@ 5.6 Reinforcement learning
Reinforcement learning (RL) has emerged as a promising approach for increasing the performance and adaptability of NER systems. Agents are trained to make sequences of decisions through the rewarding of desirable actions and the penalizing of undesirable ones. In the context of NER, RL can optimize the identification and classification of named entities by learning from interactions with data and progressively improving the model’s performance through rewards and penalties.
@@
It is also possible to combining RL with adversarial training. Adversarial training consists in training models to distinguish between true and false positives in a competitive setting, which can improve resilience to noisy and incomplete annotations. In the biomedical domain, RL improves the recognition of complex medical entities, handling a wide range of biomedical terms and variations, thus improving the extraction of meaningful entities from medical texts (Peng et al., 2021).
Studies such as (Yang et al., 2018) have shown that RL can effectively handle incomplete and noisy annotations. For instance, by integrating partial annotation learning, RL reduces the impact of unknown labels. An RL-based instance selector can filter out noisy annotations, further refining the model’s accuracy.
@@ 5.7 Graph-based methods
Further advances, particularly for Chinese NER, have led to the development of specialized graph-based networks, such as the Polymorphic Graph Attention Network (PGAT) by Wang et al. (2022c). PGAT improves character representation by dynamically integrating lexicon information. Another approach, the Lexicon-Based Graph Neural Network (LGN) by Gui et al. (2019b), constructs connections among characters, leveraging both local compositional structures and global sentence semantics. This approach demonstrated significant performance improvements across various Chinese NER datasets. Furthermore, the Multi-Graph Collaborative Network (MGCN) of Zhang et al. (2022) addresses challenges like boundary confusion and irrelevant lexical words, outperforming existing models by constructing multiple relationships between lexical words and characters in order to better capture entity boundaries.
@@ 6.1 Transfer learning
Transfer learning applies knowledge from one task in order to improve performance on a related task (Zhuang et al., 2020). It is a technique that has proven useful in areas such as image classification (Shaha and Pawar, 2018), speech recognition (Wang and Zheng, 2015), and time series classification (Fawaz et al., 2018). In NER, transfer learning is implemented by first pre-training a model on a large corpus of generic text data, followed by fine-tuning on a smaller dataset specifically designed for the target NER application. This approach utilizes the broad knowledge gained from the generic dataset so as to obtain better performance on the specific task at hand. One example of transfer learning is BERTweet, which is a BERT-based model pre-trained on Twitter data that achieves good performances on Twitter text classification, as well as on POS-tagging and named-entity recognition (Nguyen et al., 2020).
@@ 6.3 Active learning
When active learning is applied successfully to NLP, it can either improve model performance with the same amount of labeled data, or alternatively maintain similar performance while reducing the amount of data and annotation necessary. In deep learning research, active learning techniques have shown promising results. For example, Siddhant and Lipton (2018) explored the use of uncertainty-based strategies such as Least Confident (LC) and Monte Carlo Dropout Bayesian Active Learning (DO-BALD), demonstrating the ability of these techniques to reduce the annotation requirement in deep learning models (Siddhant and Lipton, 2018). Shen et al. (2017) introduced Maximum Normalized Log-Probability (MNLP), an improvement over LC that normalizes uncertainty scores by sequence length, which is particularly effective in sequence-labeling tasks like NER.
In the context of NER, several deep learning approaches have adopted active learning strategies to improve model performance with limited annotations. For example, (Yan et al., 2022) investigated approaches that combine uncertainty- and diversity-based sampling to efficiently select the most informative examples for sequence labeling. Their work illustrates how gradient embeddings and clustering techniques, such as weighted k-means++, can be used to achieve a balance between informative and diverse sample selection, thus enhancing learning efficiency in deep active learning setups (Yan et al., 2022).
@@ 6.4 Few-shot learning
In the context of named entities, studies such as Fritzler et al. (2019) and Hou et al. (2020) have proposed the adaptation of prototypical networks (Snell et al., 2017) for NER. However, those implementations were not able to achieve optimal performance. Yang and Katiyar (2020) introduced a few-shot learning method based on nearest neighbors and structured inference that is shown to be superior to classical meta-learning approaches. Cui et al. (2021) approached the NER task as a language template classification problem, outperforming traditional sequence-labeling methods. An increasing number of works are recognizing the potential of few-shot learning in NER (Hofer et al., 2018; Huang et al., 2020b; He et al., 2023).
@@ 6.5 Zero-shot learning
Zero-shot learning uses a pre-trained model to assign classes to elements that the model has never encountered previously (Larochelle et al., 2008; Lampert et al., 2013; Ding et al., 2017). This approach has been explored for linking entities (Wu et al., 2019) and typing named entities (Obeidat et al., 2019) (i.e., attributing a semantic label to a given entity).
@@ 7 Software frameworks
NER has evolved significantly, and today a number of software frameworks provide robust tools to build and deploy NER systems. These frameworks simplify the development process by offering pre-built models, customizable pipelines, and support for various languages and domains. In this section we present some of the most widely recognized and commonly used NER frameworks, highlighting their key features and capabilities.
@@
  • Apache OpenNLP (Kwartler, 2017) is a library that supports common NLP tasks such as NER, language detection, POS tagging, and chunking. Unlike other frameworks, which may use a single model for all entity types, OpenNLP provides a specialized maximum entropy model for each named entity type.
@@
  • Flair (Akbik et al., 2019) is a free, open-source library that enables the creation of NLP pipelines for multilingual applications. Flair allows the stacking of embeddings, meaning users can combine different embeddings (such as Flair, ELMo, and BERT) to improve NER performance. It supports various language models, including Flair embeddings, ELMo (Peters et al., 2018a), and BERT (Vaswani et al., 2017).
@@
  • TNER (Ushio and Camacho-Collados, 2022) is a Python library for training and tuning NER models implemented in PyTorch. It features a web application with an intuitive interface that allows users to visualize predictions.
@@
Packages like Apache OpenNLP, Stanford CoreNLP, and spaCy are also accessible in languages other than Python. For instance, openNLP [^2] https://cran.r-project.org/web/packages/openNLP/index.html is an R package that takes advantage of the capabilities of the Apache OpenNLP library, originally Java-based, by acting as an interface within the R environment. Similarly, the spacyr [^3] https://cran.r-project.org/web/packages/spacyr/index.html package connects R to spaCy. Notably, the spacyr package facilitates NER using spaCy’s pre-trained language models. Other solutions like the reticulate [^4] https://cran.r-project.org/web/packages/reticulate/index.html package make it easier to achieve interoperability between R and Python, enabling Python libraries such as Hugging Face to be accessible within R.

### v3

@@ 5.3 Deep-learning-based methods
The introduction of neural probabilistic language models by Bengio et al. (2003) laid the foundation for deep learning in NLP. By demonstrating the power of distributed word representations, Bengio’s work showed that neural networks could capture word similarities in high-dimensional spaces, a capability that would later be crucial for deep learning-based NER systems. Building on this foundation, Collobert (2011) applied deep learning to NER, proving that CNNs could be used successfully for a wide range of NLP tasks, including NER, semantic role labeling, and chunking.
@@ 5.3.1 Data representation
Word embeddings transform words into dense (or sparse) vectors of real numbers in a high-dimensional space. The resulting vectors represent each word’s semantic meaning based on surrounding words in a corpus. Traditional methods include techniques such as One-Hot encoding and TF-IDF (Ramos et al., 2003). Static word embeddings, such as Word2Vec (Mikolov, 2013), GloVe (Pennington et al., 2014), and fastText (Bojanowski et al., 2016), provide a fixed vector for each word regardless of its context. Contextual embeddings, based on transformers such as GPT (Radford, 2018) and BERT (Devlin et al., 2018), capture the meaning of words based on their specific context within a sentence, allowing for more nuanced and flexible word representations.
@@
Modern word embeddings like Word2Vec and GloVe produce dense vector representations, where similar terms have close vectors. These embeddings are derived from large corpora in an unsupervised manner. Word2Vec uses Continuous Bag-of-Words (CBOW), which predicts a target word from its context, and Skip-gram, which predicts context words from a target word. Both architectures are trained with noise contrastive estimation (NCE), a type of negative sampling. NCE increases the likelihood of the target word in context and decreases the likelihood of noise words, effectively teaching the model to distinguish between true contexts and artificially generated noise (Mikolov et al., 2013). This contrastive approach leads to dense embeddings that are highly effective for capturing semantic similarities, as it forces the model to focus on discriminative features during training, enhancing the quality of the resulting vector representations. GloVe, on the other hand, uses matrix factorization of the word co-occurrence matrix, capturing the corpus’s general statistics for efficient, high-performance embeddings (Pennington et al., 2014). Unlike Word2Vec, which is context-based, GloVe emphasizes global word co-occurrence to produce vectors that represent the global relationships between words.
Numerous studies, including (Collobert et al., 2011; Huang et al., 2015), have used word embeddings for NER. For example, Ma and Hovy (2016) evaluated the performance of their NER system using different word embeddings such as Word2Vec and GloVe. The importance of these embeddings to achieve good performance is presented in (Lample, 2016).
Finally, we note that word embeddings can also be combined, as in (Dadas, 2019; Das et al., 2017), where a Wikipedia knowledge base is used to annotate named entities. In (Dadas, 2019), the labels are transformed into One-Hot vectors and concatenated with Word2Vec or ELMo (Peters et al., 2018) word embeddings.
Character embeddings represent words as sequences of character vectors, capturing their internal structure. Unlike traditional embeddings, they vectorize characters and combine them to form word representations, as Figure 6 illustrates in the case of a CNN architecture.
@@
Character embeddings can be created using several methods. One-hot encoding is a technique that represents characters as binary vectors with a 1 at the index corresponding to the character. Chars2Vec [^1] https://github.com/IntuitionEngineeringTeam/chars2vec?tab=readme-ov-file, inspired by Word2Vec, generates embeddings by predicting characters based on their neighboring characters. CNN-based embeddings treat character sequences as images, using CNN filters to extract meaningful features. BiLSTM-based embeddings use bidirectional Long Short-Term Memory (Hochreiter, 1997; Graves and Graves, 2012) networks to capture contextual information from character sequences.
CNN and RNN, which we describe in the following section, are used in (Ma and Hovy, 2016; Peters et al., 2017; Zhou et al., 2023a) to calculate feature vectors for each word. One of the conclusions drawn in (Lample, 2016) is that recurrent models tend to prioritize later elements, leading to feature vectors that represent suffixes more strongly than prefixes. Consequently, the authors suggest the use of BiLSTM to more effectively capture the prefix information.
@@ 5.3.2 Context encoding
It is for this reason that Huang et al. (2015) introduced an LSTM model for NER and showed that adding a CRF layer as a tag decoder can improve performance. This combination of LSTM for sequence modeling and CRF for decoding proved to be a highly effective strategy. In other areas, similar approaches were used by Chalapathy et al. (2016) for Drug NER and Zhang and Yang (2018) for Chinese NER. An RNN based on a BiLSTM framework was used in (Huang et al., 2015); the same approach has since been adopted by other authors (Ma and Hovy, 2016; Lample, 2016; Huang et al., 2020a). This widespread adoption underscores the robustness and flexibility of RNN-based models for a variety of NER applications.
@@ 5.3.3 Entity decoding
CRF, for example, is a probabilistic model that assigns labels by considering the entire sequence of tokens rather than making independent predictions for each one. It excels in modeling dependencies between adjacent labels, which is crucial for ensuring coherent tag sequences. By learning to assign higher probabilities to valid label sequences (such as recognizing that a "Person" label is likely to be followed by a "Location"), CRF provides a robust method for NER. Studies by Lample (2016) and Ma and Hovy (2016) show that incorporating a CRF layer into deep learning models, such as BiLSTM (Luo et al., 2018; Lin et al., 2019) or CNN (Knobelreiter et al., 2017; Feng et al., 2020), enhances performance. This combination leverages CRF’s ability to ensure consistent sequence labeling while benefiting from the feature extraction power of neural networks. Many advanced NER models integrate a CRF layer with either BiLSTM or CNN for feature extraction. The BiLSTM captures the sequential context by processing the input in both forward and backward directions, ensuring that the model considers the entire sentence before making predictions. CNN, on the other hand, focuses on local patterns such as prefixes, suffixes, and word combinations, which are key in identifying entities. The CRF layer is then used to refine the final output, ensuring consistency across the sequence. This combination allows the model to capture both local features and long-range dependencies, making it more accurate and reliable.
@@ 5.4 Transformer-based language methods
Although Transformer-based models such as BERT, RoBERTa, DeBERTa, and GLiNER have shown excellent performances in NER, they are computationally intensive. The motivation behind attempts to create more efficient models such as DistilBERT is obtaining a better balance between performance and resource requirements. Recent advances also include the development of multilingual models such as mBERT (Devlin et al., 2018), which can handle multiple languages and be fine-tuned for specific languages using bilingual gazetteers or additional datasets, enhancing performance in low-resource settings. Additionally, models such as Language Understanding with Knowledge-based Embeddings (LUKE) incorporate entity information into word embeddings, improving entity recognition by using innovative masking and self-attention techniques (Yamada et al., 2020).
@@ 5.5.1 Principles and applications of LLMs in NER
LLMs are based on the Transformer decoder architecture, in which a multitude of attention mechanisms are orchestrated in layers to form an intricate neural network. The structural designs and pre-training paradigms implemented in current LLMs exhibit strong parallels with those used in smaller-scale language models. The primary distinction is the markedly increased size of both the model parameters and the training corpus. Some LLMs, such as T5 (Raffel et al., 2020), operate as hybrids, incorporating the encoder and decoder modules of the Transformer to increase comprehension and generative functionalities.
@@ 5.5.2 Pros and Cons of LLMs for NER
Another important advantage is contextual understanding. The transformer-based architecture of LLMs enables these models to capture complex contextual relationships, an essential feature for NER tasks in specialized domains, such as medical or historical texts. For example, models such as Resolving ECCNPs (Kammer et al., 2023), customized for German medical texts, effectively handle unique language constructs and domain-specific terminology. This ability to understand nuanced context improves LLM performance in scenarios where accurate interpretation of specialized language is crucial.
@@ 5.6 Reinforcement learning
RL has emerged as a promising approach for increasing the performance and adaptability of NER systems. Agents are trained to make sequences of decisions through the rewarding of desirable actions and the penalizing of undesirable ones. In the context of NER, RL can optimize the identification and classification of named entities by learning from interactions with data and progressively improving the model’s performance through rewards and penalties.
@@
It is also possible to combine RL with adversarial training. Adversarial training consists in training models to distinguish between true and false positives in a competitive setting, which can improve resilience to noisy and incomplete annotations. In the biomedical domain, RL improves the recognition of complex medical entities, handling a wide range of biomedical terms and variations, thus improving the extraction of meaningful entities from medical texts (Peng et al., 2021).
Studies such as (Yang et al., 2018) have shown that RL can effectively handle incomplete and noisy annotations. For example, by integrating partial annotation learning, RL reduces the impact of unknown labels. An RL-based instance selector can filter out noisy annotations, further refining the model’s accuracy.
@@ 5.7 Graph-based methods
Further advances, particularly for Chinese NER, have led to the development of specialized graph-based networks, such as the Polymorphic Graph Attention Network (PGAT) by Wang et al. (2022b). PGAT improves character representation by dynamically integrating lexicon information. Another approach, the Lexicon-Based Graph Neural Network (LGN) by Gui et al. (2019b), constructs connections among characters, leveraging both local compositional structures and global sentence semantics. This approach demonstrated significant performance improvements across various Chinese NER datasets. Furthermore, the Multi-Graph Collaborative Network (MGCN) of Zhang et al. (2022) addresses challenges like boundary confusion and irrelevant lexical words, outperforming existing models by constructing multiple relationships between lexical words and characters in order to better capture entity boundaries.
@@ 6.1 Transfer learning
Transfer learning applies knowledge from one task in order to improve performance on a related task. It is a technique that has proven useful in areas such as image classification (Shaha and Pawar, 2018), speech recognition (Wang and Zheng, 2015), and time series classification (Fawaz et al., 2018). In NER, transfer learning is implemented by first pre-training a model on a large corpus of generic text data, followed by fine-tuning on a smaller dataset specifically designed for the target NER application. This approach utilizes the broad knowledge gained from the generic dataset so as to obtain better performance on the specific task at hand. One example of transfer learning is BERTweet, which is a BERT-based model pre-trained on Twitter data that achieves good performances on Twitter text classification, as well as on POS-tagging and named-entity recognition (Nguyen et al., 2020).
@@ 6.3 Active learning
When active learning is applied successfully to NLP, it can either improve model performance with the same amount of labeled data, or alternatively maintain similar performance while reducing the amount of data and annotation necessary. In deep learning research, active learning techniques have shown promising results. For example, Siddhant and Lipton (2018) explored the use of uncertainty-based strategies such as Least Confident (LC) and Monte Carlo Dropout Bayesian Active Learning (DO-BALD), demonstrating the ability of these techniques to reduce the annotation requirement in deep learning models. Shen et al. (2017) introduced Maximum Normalized Log-Probability (MNLP), an improvement over LC that normalizes uncertainty scores by sequence length, which is particularly effective in sequence-labeling tasks like NER.
In the context of NER, several deep learning approaches have adopted active learning strategies to improve model performance with limited annotations. For example, (Yan et al., 2023) investigated approaches that combine uncertainty- and diversity-based sampling to efficiently select the most informative examples for sequence labeling. Their work illustrates how gradient embeddings and clustering techniques, such as weighted k-means++, can be used to achieve a balance between informative and diverse sample selection, thus enhancing learning efficiency in deep active learning setups.
@@ 6.4 Few-shot learning
In the context of named entities, studies such as Fritzler et al. (2019) and Hou et al. (2020) have proposed the adaptation of prototypical networks (Snell et al., 2017) for NER. However, these implementations were unable to achieve optimal performance. Yang and Katiyar (2020) introduced a few-shot learning method based on nearest neighbors and structured inference that is shown to be superior to classical meta-learning approaches. Cui et al. (2021) approached the NER task as a language template classification problem, outperforming traditional sequence-labeling methods. An increasing number of works are recognizing the potential of few-shot learning in NER (Hofer et al., 2018; Huang et al., 2020b; He et al., 2023).
@@ 6.5 Zero-shot learning
Zero-shot learning uses a pre-trained model to assign classes to elements that the model has never encountered previously (Larochelle et al., 2008; Lampert et al., 2013; Ding et al., 2017). This approach has been explored for linking entities (Wu et al., 2020) and typing named entities (Obeidat et al., 2019) (i.e., attributing a semantic label to a given entity).
@@ 7 Software frameworks
NER has evolved significantly, and today a number of software frameworks provide robust tools to build and deploy NER systems. These frameworks simplify the development process by offering pre-built models, customizable pipelines, and support for various languages and domains. In this section, we present some of the most widely recognized and commonly used NER frameworks, highlighting their key features and capabilities.
@@
  • Apache OpenNLP (Baldridge, 2005) is a library that supports common NLP tasks such as NER, language detection, POS tagging, and chunking. Unlike other frameworks, which may use a single model for all entity types, OpenNLP provides a specialized maximum entropy model for each named entity type.
@@
  • Flair (Akbik et al., 2019) is a free, open-source library that enables the creation of NLP pipelines for multilingual applications. Flair allows the stacking of embeddings, meaning users can combine different embeddings (such as Flair, ELMo, and BERT) to improve NER performance. It supports various language models, including Flair embeddings, ELMo, and BERT.
@@
  • TNER (Ushio and Camacho-Collados, 2021) is a Python library for training and tuning NER models implemented in PyTorch. It features a web application with an intuitive interface that allows users to visualize predictions.
@@
Packages like Apache OpenNLP, Stanford CoreNLP, and spaCy are also accessible in languages other than Python. For example, openNLP [^2] https://cran.r-project.org/web/packages/openNLP/index.html is an R package that takes advantage of the capabilities of the Apache OpenNLP library, originally Java-based, by acting as an interface within the R environment. Similarly, the spacyr [^3] https://cran.r-project.org/web/packages/spacyr/index.html package connects R to spaCy. Notably, the spacyr package facilitates NER using spaCy’s pre-trained language models. Other solutions like the reticulate [^4] https://cran.r-project.org/web/packages/reticulate/index.html package make it easier to achieve interoperability between R and Python, enabling Python libraries such as Hugging Face to be accessible within R.